OpenStudy (twistrose):

A clerk takes 3/4 h to process 9 applications. How long does it take to process 1 application?

OpenStudy (igreen):

You can set up a proportion. \(\sf \dfrac{0.75}{9} = \dfrac{x}{1}\) Cross multiply, can you do that?

OpenStudy (vijeya3):

For 9 hours it is 3/4..therefore for 1 hour it would be (3/4)?9 =1/12 hours (approx)0.083 hours If this is helpful,medal me:)

OpenStudy (twistrose):

thanks :)